# Near-Optimal Regret for Efficient Stochastic Combinatorial Semi-Bandits

### v1

@@ 3.2 Regret analysis for semi-bandit feedback
In this case $r(A_{t},\bm{\mu})=\sum_{i\in A_{t}}\mu_{i}$. Since $\mathcal{A}$ consists of all subsets of $[m]$ with cardinality less than $k$, and the reward function is monotonically increasing, then $A_{t}$ is the set of $k$ arms corresponding to the largest values of $\bar{\mu}_{i}$. Since $\Delta_{A_{t}}=\sum_{j\in A^{*}}\mu_{j}-\sum_{i\in A_{t}}\mu_{i}$, we have
@@
    Decoupling constraints across suboptimal action $A_{t}$ into individual arms. We classify the action set into well-defined categories (Lemma 2) and quantitatively characterize the necessary condition that must be satisfied by the suboptimal arms belong to $A_{t}$ in each category (Lemma 3);
@@
All proofs of the lemmas in this section are deferred to Appendix A.1. Here $\delta$ is optimized to be $\frac{m}{kT}$. Define $\zeta_{j}=\max\left(0,\mu_{j}-\min\limits_{t\leq T}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\right)$ for every $j\in A^{*}$. We decompose the overall regret as follows:
@@
$\leq\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq\sqrt{\frac{ekm}{T}}\right]\Delta_{A_{t}}\right]$
$\phantom{{}={}}+\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq 8\sum_{j\in A^{*}}\zeta_{j}\right]\Delta_{A_{t}}\right]$
$\phantom{{}={}}+\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$\leq\sqrt{ekmT}+8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T$
$\phantom{{}={}}+\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right].$ (1)
From Lemma 8 in Appendix A.1, which is proofed by using concentration inequalities, the second term could be bounded as
$8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T\leq 32k\sqrt{\delta}\cdot T=32\sqrt{kmT}.$
@@
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right\}.$
@@
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\log k\sqrt{kmT}\right).$
@@
$=O\left((\log k)^{2}\sqrt{kmT}\right).$
Bring this back to eq. 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)^{2}\sqrt{kmT}\right)$.
@@ 3.3 Extension to cascading feedback
In this section we introduce the cascading feedback (Kveton et al. 2015a, b; Li et al. 2016) and give our regret upper bound. In this model, when $A_{t}$ is chosen, playing this action means that the player reveals the outcomes of the arms one by one following the sequence order until certain stopping condition is satisfied. The feedback is the outcomes of revealed arms and the reward is a function form of all arms in $A_{t}$. In particular, in the disjunctive form the player stops when the first 1 is revealed, or he reaches the end and gets all feedbacks. In the conjunctive form, the player stops when the first 0 is revealed or she reaches the end with all 1 outcomes.
@@
For any stochastic combinatorial cascading feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then the expected regret of Algorithm 1 is $O\left(\frac{(\log k)^{2}}{p^{*}}\sqrt{kmT}\right)$.
The proof is given in Appendix A.2. In Appendix C.4, we also present comparative experiments evaluating the performance of CUCB and CMOSS under cascading feedback.
@@ 4 Experiments
Figure 1: Comparison of CMOSS (blue) with three baselines algorithms under semi-bandit feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=10$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.5,0.6]$. Shaded regions and error bars denote one standard deviation over 10 runs.
@@
For all experiments, we fix the confidence parameter of CMOSS to $\delta=0.00001$, which is of the same order of magnitude as the theoretical result $\frac{m}{kT}$ (see Section 3.1). For EXP3.M, we use a mixing coefficient $\gamma=0.01$, as it yields favorable empirical performance in our settings (see Appendix C.3). All rewards follow Bernoulli distributions: pulling arm $\mathtt{arm}_{i}$ yields reward $1$ with probability $\mu_{i}$, and 0 otherwise, where $\mu_{i}$ denotes the true mean (interpretable as the latent click-through probability in recommendation tasks).
@@ 4.1 Comparison with baselines
We evaluate CMOSS across both synthetic and real-world bandit environments to assess its effectiveness under diverse reward structures. (1)(2)(5)(6) of Figure 1 report the cumulative regret over time with fixed cardinality $k=10$ and number of base arms $m=30$.
In the synthetic setting, the initial click probability $\mu_{i}$ for each arm $\mathtt{arm}_{i}$ is independently drawn from a uniform distribution. We consider two regimes: a low-probability regime with $\mu_{i}\sim\mathtt{U}(0,0.1)$, and a high-probability regime with $\mu_{i}\sim\mathtt{U}(0.5,0.6)$. To further assess the practical effectiveness of our method, we evaluate it on the public Yelp dataset [^1] http://www.yelp.com/dataset˙challenge, a classical CMAB real-world dataset. Following (Li, Chen, and Leung 2019), we map user–item affinities to the target regimes $[0,0.1]$ and $[0.5,0.6]$ via linear rescaling. Full details are provided in Appendix C.1. The resulting values serve as the initial click probabilities of base arms in the two settings. For each regime, we independently sample $30$ arms using a fixed random seed ($2025$) to ensure reproducibility.
The low-probability regime $[0,0.1]$ reflects typical conditions observed in real-world recommendation scenarios, where user interactions such as clicks occur relatively infrequently. This setting provides a realistic environment for evaluating algorithm performance. In contrast, the high-probability regime $[0.5,0.6]$ represents a more idealized scenario with higher click probabilities, allowing us to assess algorithm behavior in more optimistic conditions. By examining both regimes, we gain complementary insights into the robustness and effectiveness of the algorithms across a range of practical and theoretical contexts.
@@ 4.2 Ablation studies
To further evaluate the robustness and scalability of CMOSS, we conduct ablation studies by varying cardinality $k$ while fixing base arm size at $m=30$ and varying base arm size $m$ while fixing cardinality $k=10$ on synthetic dataset. As shown (3)(4) (varyin $k$) and (7)(8) (varyin $m$) of Figure 1, we report both cumulative regret and total runtime. Detailed metrics are summarized in Table 3. We also do the same ablation studies on Yelp dataset (see details in Appendix C.2).
As demonstrated, CMOSS consistently achieves the lowest cumulative regret across varying values of $k$ and $m$, outperforming all baseline algorithms. Compared to CUCB, CMOSS attains more than a 45% reduction in regret while typically incurring less than a 20% increase in runtime. Moreover, relative to EXP3.M and HYBRID, CMOSS not only achieves 25% to 70% lower regret but also reduces runtime by over 75%. The ablation studies on the Yelp dataset exhibit the same trends, further confirming the robustness and efficiency of CMOSS across different data domains.
In addition to the main experiments, we include supplementary studies in Appendix C to strengthen the completeness and rigor of our evaluation. We first perform a parameter selection study for EXP3.M (see Appendix C.3) to validate the choice of its hyperparameter $\gamma=0.01$ in our setup. We then compare CMOSS and CUCB under a cascading feedback model (see Appendix C.4) to assess their performance under an alternative user interaction mechanism. In particular, CMOSS achieves a notable reduction in regret under cascading feedback, while incurring only a modest increase in runtime. These additional experiments support both the soundness of our experimental design and the robustness of CMOSS across varied settings.
@@ 5 Conclusion and Future Work
In this paper, we study the stochastic CMAB problem and propose the CMOSS algorithm, which is not only computationally efficient but also has a regret of $O\left((\log k)^{2}\sqrt{kmT}\right)$ under semi-bandit feedback. We then extend our analysis to cascading feedback. Comparative experiments with CUCB, EXP3.M and HYBRID under semi-bandit feedback show that CMOSS has at least one-way significant advantages in cumulative regret and computation cost over the other three algorithms, with the other one showing small differences.
@@ Appendix A Techinical Proofs of Section 3
For completeness and consistency of the paper, we first provide proofs of some basic lemmas (Lemma 7, 8, 9) that will be used later.
@@
where eq. 2 follows from Lemma 7. And eq. 3 follows from
@@
Then we begin to give a thorough proof of Theorem 1. We focus on proving the third term $\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$, where
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right\}.$
When Algorithm 1 chooses $A_{t}$, for the arms in $A_{t}\setminus A^{*}$ and arms in $A^{*}\setminus A_{t}$, we have
$\min\limits_{i\in A_{t}\setminus A^{*}}\left(\hat{\mu}_{i,t}+\rho_{i,t}\right)\geq\max\limits_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\sum_{j\in A_{t}\setminus A^{*}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right).$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\min\limits_{s\leq T}\left(\sum_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right)\right)\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\sum_{j\in A^{*}\setminus A_{t}}\min\limits_{s\leq T}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right).$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}$ $\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\sum_{j\in A^{*}\setminus A_{t}}\left(\mu_{j}-\zeta_{j}\right)\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)\geq\frac{1}{\left|A_{t}\setminus A^{*}\right|}\left(\sum_{i\in A_{t}\setminus A^{*}}\mu_{i}+\frac{7\Delta_{A_{t}}}{8}\right).$
Remember that we denote $\left(\mu^{A}_{i}\right)_{i\in[k_{A}]}$ as the mean rewards of the arms in $A\setminus A^{*}$ where $\mu^{A}_{k_{A}}\geq\mu^{A}_{k_{A}-1}\geq\cdots\geq\mu^{A}_{1}$ and $\overline{\mu^{A}}=\frac{\sum_{i\in[k_{A}]}\mu^{A}_{i}}{k_{A}}$. Define $D_{A}=\left\{i\in[k_{A}]\;\middle|\;\mu^{A}_{i}\leq\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}\right\}$ as the set of arms in $A\setminus A^{*}$ whose mean reward is no more than the $\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}$. Then we have Lemma 2:
@@
which contradicts eq. 4. ∎
From Lemma 2, before the algorithm starts, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log k\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log k\rfloor}$, where
@@
$\phantom{{}={}}\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\land\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\leq\frac{2}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\Delta_{A_{t}}}{16k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]$ $\leq\frac{1}{\lfloor\frac{k_{A_{t}}}{2^{\ell}}\rfloor}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right]$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\land\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\land\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\log k\sqrt{kmT}\right).$
@@
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\land\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$

### v2

@@ 3.2 Regret analysis for semi-bandit feedback
In this case $r(A_{t},\bm{\mu})=\sum_{i\in A_{t}}\mu_{i}$. Since $\mathcal{A}$ consists of all subsets of $[m]$ with cardinality less than $k$, and the reward function is monotonically increasing, then $A^{*}$ consists of $k$ arms corresponding to the largest values of $\mu_{i}$. $A_{t}$ is the set of $k$ arms corresponding to the largest values of $\bar{\mu}_{i}$ in round $t$. Since $\Delta_{A_{t}}=\sum_{j\in A^{*}}\mu_{j}-\sum_{i\in A_{t}}\mu_{i}$, we have
@@
    Decoupling constraints across suboptimal action $A_{t}$ into individual arms. We classify the action set into well-defined categories (Lemma 2) and quantitatively characterize the necessary condition that must be satisfied by the suboptimal arms belong to $A_{t}$ in each category Lemma 3);
@@
In this section we only present the case where $k\leq\frac{m}{2}$. All proofs of the lemmas in this section and the other case are deferred to Section A.1. Here $\delta$ is optimized to be $\frac{m(\log k)^{2}}{kT}$. Define $\zeta_{j}=\max\left(0,\mu_{j}-\min\limits_{t\leq T}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\right)$ for every $j\in A^{*}$. We decompose the overall regret as follows:
@@
$\leq\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq(\log k)\sqrt{\frac{ekm}{T}}\right]\Delta_{A_{t}}\right]+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}\leq 8\sum_{j\in A^{*}}\zeta_{j}\right]\Delta_{A_{t}}\right]+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$\leq\sqrt{ekmT}+8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T+$
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right].$ (1)
From Lemma 8 in Section A.1, which is proved by using concentration inequalities, the second term could be bounded as
$8\sum_{j\in A^{*}}\mathbf{E}\left[\zeta_{j}\right]\cdot T\leq 32k\sqrt{\delta}\cdot T=32(\log k)\sqrt{kmT}.$
@@
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right\}.$
@@
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\sqrt{kmT}\right).$
@@
$=O\left((\log k)\sqrt{kmT}\right).$
Bring this back to Equation 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)\sqrt{kmT}\right)$.
@@ 3.3 Extension to cascading feedback
In this section we introduce the cascading feedback (Kveton et al. 2015a, b; Li et al. 2016) and give our regret upper bound. In this model, when $A_{t}$ is chosen, playing this action means that the player reveals the outcomes of arms one by one following the sequence order until certain stopping condition is satisfied. The feedback is the outcomes of revealed arms and the reward is a function form of all arms in $A_{t}$. In particular, in the disjunctive form the player stops when the first 1 is revealed, or he reaches the end and gets all feedback. In the conjunctive form, the player stops when the first 0 is revealed or she reaches the end with all 1 outcomes.
@@
For any stochastic combinatorial cascading feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then the expected regret of Algorithm 1 is $O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)$ when $k>\frac{m}{2}$.
The proof is given in Section A.2. In Section C.4, we also present comparative experiments evaluating the performance of CUCB and CMOSS under cascading feedback.
@@ 4 Experiments
Figure 1: Comparison of CMOSS (blue) with three baselines algorithms under semi-bandit feedback. Subplots (1)-(4), (9)-(12) use synthetic dataset; (5)-(8) use Yelp dataset. Subplots (1)-(8) show cumulative regret with $k=10,m=30$ and $k=20,m=30$; (9)(10) show ablation studies varying $k$ (fixed $m=30$), while (11)(12) varying $m$ (fixed $k=15$). Initial means of base arms fall within the range $[0,0.1]$, except for (2)(4)(6)(8), which use the range $[0.3,0.4]$.
@@
For all experiments, we fix the confidence parameter of CMOSS to $\delta=0.00001$, which is of the same order of magnitude as the theoretical results $\frac{m(\log k)^{2}}{kT},k\leq\frac{m}{2}$ and $\frac{(m-k)^{2}\log k\log(m-k)}{k^{2}T},k>\frac{m}{2}$ (see Section 3.1 and Section A.1). For EXP3.M, we use a mixing coefficient $\gamma=0.01$, as it yields favorable empirical performance in our settings (see Section C.3). All rewards follow Bernoulli distributions: pulling arm $\mathtt{arm}_{i}$ yields reward $1$ with probability $\mu_{i}$, and 0 otherwise, where $\mu_{i}$ denotes the true mean (interpretable as the latent click-through probability in recommendation tasks).
@@ 4.1 Comparison with baselines
We evaluate CMOSS across both synthetic and real-world bandit environments to assess its effectiveness under diverse reward structures. (1)-(8) of Figure 1 report the cumulative regret over time with cardinality $k=10$ and $k=20$ and fixed number of base arms $m=30$.
In the synthetic setting, the initial click probability $\mu_{i}$ for each arm $\mathtt{arm}_{i}$ is independently drawn from a uniform distribution. We consider two regimes: a low-probability regime with $\mu_{i}\sim\mathtt{U}(0,0.1)$, and a high-probability regime with $\mu_{i}\sim\mathtt{U}(0.3,0.4)$. To further assess the practical effectiveness of our method, we evaluate it on the public Yelp dataset [^1] http://www.yelp.com/dataset˙challenge, a classical CMAB real-world dataset. Following (Li, Chen, and Leung 2019), we map user–item affinities to the target regimes $[0,0.1]$ and $[0.3,0.4]$ via linear rescaling. Full details are provided in Section C.1. The resulting values serve as the initial click probabilities of base arms in the two settings. For each regime, we independently sample $30$ arms.
The low-probability regime $[0,0.1]$ reflects typical conditions observed in real-world recommendation scenarios, where user interactions such as clicks occur relatively infrequently. This setting provides a realistic environment for evaluating algorithm performance. In contrast, the high-probability regime $[0.3,0.4]$ represents a more idealized scenario with higher click probabilities, allowing us to assess algorithm behavior in more optimistic conditions. By examining both regimes, we gain complementary insights into the robustness and effectiveness of the algorithms across a range of practical and theoretical contexts.
@@ 4.2 Ablation studies
To further evaluate the robustness and scalability of CMOSS, we conduct ablation studies by varying cardinality $k$ while fixing base arm size at $m=30$ and varying base arm size $m$ while fixing cardinality $k=15$ on synthetic dataset. As shown (3)(4) (varying $k$) and (7)(8) (varying $m$) of Figure 1, we report both cumulative regret and total runtime. Detailed metrics are summarized in Table 3. We also do the same ablation studies on Yelp dataset (see details in Section C.2).
Across both synthetic and Yelp datasets, CMOSS consistently achieves the lowest cumulative regret when varying either $k$ or $m$. Relative to CUCB, CMOSS reduces regret by 40–55% in most configurations, while requiring only a 10–30% increase in runtime. When compared with EXP3.M and HYBRID, CMOSS offers significant improvements: regret is typically 25–70% lower, and runtime is reduced by more than 65–85% across nearly all settings. The ablation studies on the Yelp dataset reveal the same performance patterns as those on the synthetic dataset, confirming that CMOSS remains robust and efficient across different domains and parameter regimes.
In addition to the main experiments, we include supplementary studies in Appendix C to strengthen the completeness and rigor of our evaluation. We first perform a parameter selection study for EXP3.M (see Section C.3) to validate the choice of its hyperparameter $\gamma=0.01$ in our setup. We then compare CMOSS and CUCB under a cascading feedback model (see Section C.4) to assess their performance under an alternative user interaction mechanism. In particular, CMOSS achieves a notable reduction in regret under cascading feedback, while incurring only a modest increase in runtime. These additional experiments support both the soundness of our experimental design and the robustness of CMOSS across varied settings.
@@ 5 Conclusion and Future Work
In this paper, we study the stochastic CMAB problem and propose the CMOSS algorithm, which is not only computationally efficient but also has a regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$ under semi-bandit feedback. We then extend our analysis to cascading feedback. Comparative experiments with CUCB, EXP3.M and HYBRID under semi-bandit feedback show that CMOSS has at least one-way significant advantages in cumulative regret and computation cost over the other three algorithms, with the other one showing small differences.
@@ Appendix A Techinical Proofs of Section 3
For completeness and consistency of the paper, we first provide proofs of some basic lemmas (Lemma 7, Lemma 8, Lemma 9) that will be used later.
@@
where Equation 2 follows from Lemma 7. And Equation 3 follows from
@@
We begin to give a thorough proof of Theorem 1. When Algorithm 1 chooses $A_{t}$, for the arms in $A_{t}\setminus A^{*}$ and arms in $A^{*}\setminus A_{t}$, we have
$\min\limits_{i\in A_{t}\setminus A^{*}}\left(\hat{\mu}_{i,t}+\rho_{i,t}\right)\geq\max\limits_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,t}+\rho_{j,t}\right).$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\min\limits_{s\leq T}\left(\sum_{j\in A^{*}\setminus A_{t}}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right)\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\min\limits_{s\leq T}\left(\hat{\mu}_{j,s}+\rho_{j,s}\right).$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}$ $\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\sum_{j\in A^{*}\setminus A_{t}}\left(\mu_{j}-\zeta_{j}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)$
@@
$\hat{\mu}_{i,t}+\rho_{i,t}\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{j\in A^{*}\setminus A_{t}}\mu_{j}-\sum_{j\in A^{*}}\zeta_{j}\right)\geq\frac{1}{\left|A^{*}\setminus A_{t}\right|}\left(\sum_{i\in A_{t}\setminus A^{*}}\mu_{i}+\frac{7\Delta_{A_{t}}}{8}\right).$
Remember that we denote $\left(\mu^{A}_{i}\right)_{i\in[k_{A}]}$ as the mean rewards of the arms in $A\setminus A^{*}$ where $\mu^{A}_{k_{A}}\geq\mu^{A}_{k_{A}-1}\geq\cdots\geq\mu^{A}_{1}$ and $\overline{\mu^{A}}=\frac{\sum_{i\in[k_{A}]}\mu^{A}_{i}}{k_{A}}$. Define $D_{A}=\left\{i\in[k_{A}]\;\middle|\;\mu^{A}_{i}\leq\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}\right\}$ as the set of arms in $A\setminus A^{*}$ whose mean reward is no more than $\overline{\mu^{A}}+\frac{3\Delta_{A}}{4k_{A}}$. Then we have Lemma 2:
@@
which contradicts Equation 4. ∎
Now we divide the proof into two cases where $k\leq\frac{m}{2}$ and $k>\frac{m}{2}$.
The case where $k\leq\frac{m}{2}$
We focus on proving the third term $\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$, where
$\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},(\log k)\sqrt{\frac{ekm}{T}}\right\}\right\}.$
From Lemma 2, before the algorithm starts, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log k\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log k\rfloor}$ since $1\leq k_{A}\leq k$, where
@@
$\phantom{{}={}}\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{0}\cap\mathcal{A}_{B}\right]\leq\frac{2}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\Delta_{A_{t}}}{16k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]$ $\leq\frac{1}{\lfloor\frac{k_{A_{t}}}{2^{\ell}}\rfloor}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right]$
@@
$\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\mathbb{1}\left[\left(\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\right)\land i\in A_{t}\right].$
@@
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left(\sqrt{kmT}\right).$
@@
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$